OpenStudy (anonymous):

If f(x1)+f(x2)=f(x1+x2) for all real numbers x1 and x2, which of the following could define f? A) f(x)=x+1 B)f(x)=2x C)f(x)=1/x D)f(x)=e^x E)f(x)=x^2 I have no idea how to go about this, if someone can explain it to me.

hartnn (hartnn):

given f(x), if you want to find f(x1), replace all 'x' in f(x) by 'x1' for example, if f(x) = x^3 - x f(x1) = (x1)^3 - x1 got this ?

OpenStudy (anonymous):

No I do not.

hartnn (hartnn):

which part, u have doubt with ?

OpenStudy (anonymous):

The whole thing itself. The wording is confusing.

hartnn (hartnn):

what you have to do is to find, for each option (1) f(x1)+f(x2) (2) f(x1+x2) and check, for which option are these two equal.

hartnn (hartnn):

if, x1 , x2 confuses you, you can find (1) f(a)+f(b) (2) f(a+b) where f(a) is found by replacing 'x' in f(x) by a. f(b) is found by replacing 'x' in f(x) by b. and check whether they come out to be equal.

hartnn (hartnn):

Let me do option A) for you, you check for others. f(x1) = x1+1 f(x2) = x2+1 f(x1+x2) = x1+x2+1 but f(x1)+f(x2) = x1+1 +x2+1 = x1 + x2+2 which does not equal f(x1+x2). so, option A is not correct. try others.

OpenStudy (anonymous):

I get it now. Thanks!

hartnn (hartnn):

which option you got ?

OpenStudy (anonymous):

I ended up with E.

hartnn (hartnn):

well, E is incorrect. f(x1) = x1^2 f(x2) = x2^2 f(x1+x2) = (x1+x2)^2 = x1^2+2x1*x2 + x2^2 but f(x1)+f(x2) = x1^2+x2^2 only which does not equal f(x1+x2). so, option E is not correct.

hartnn (hartnn):

did u try B) ??

OpenStudy (anonymous):

Yes. I ended up with D.

hartnn (hartnn):

even, thats incorrect....try B)

hartnn (hartnn):

were you just guessing ? show your work with option B)

OpenStudy (anonymous):

I was not guessing. Like I said, I have no idea what I'm doing.

hartnn (hartnn):

i gave you an idea with that explanation..why don't you follow that for option B)